OpenStudy (anonymous):

8000 liters to cubic liters

OpenStudy (anonymous):

8

OpenStudy (anonymous):

There's actually a question to it too

OpenStudy (stuck_in_the_middle):

8

OpenStudy (anonymous):

wat is the question?

OpenStudy (anonymous):

A 20 liter sample of gas is kept under a constant pressure of 100 kPa. The temp is lowered from 360 K to 90 K. What is the new volume

OpenStudy (anonymous):

should've been the main question!

OpenStudy (anonymous):

Ha so would it still be 8 cubic liters?

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

Use gas laws, V1/T1 = V2/T2

OpenStudy (anonymous):

guide me through? not sure where to palce the numbers.

OpenStudy (anonymous):

this is the math part of physics where I'm bad again. :P Proves my calc too

OpenStudy (anonymous):

wait let me try

OpenStudy (anonymous):

100/20 = 360/90 5/4 = 1.25 is that it?

OpenStudy (agent0smith):

idk, i'll check in a sec, just plug the numbers in where they belong. Initial and finals

OpenStudy (anonymous):

still can't get it. wha do you think?

OpenStudy (agent0smith):

Identify what all the variables are. What is T1, V1 etc

OpenStudy (anonymous):

V1 = 100 Kpa t1= 360k t2 = 90k

OpenStudy (anonymous):

I don't think i'm right. ha physics.

OpenStudy (agent0smith):

KPa is a unit of pressure.

OpenStudy (anonymous):

V1 = 360k v2 = 90k t1 = 100 kpa?

OpenStudy (anonymous):

Is that right?

OpenStudy (agent0smith):

KPa is a unit of pressure.

OpenStudy (agent0smith):

uh wow you're all over the place :P V1 = 360k v2 = 90k V1 and V2 are volumes, not temperatures.

OpenStudy (agent0smith):

A 20 liter sample of gas is kept under a constant pressure of 100 kPa. The temp is lowered from 360 K to 90 K. What is the new volume read carefully and identify each quantity that you'll need in V1/T1 = V2/T2

OpenStudy (anonymous):

so 100 kpa is the t1?

OpenStudy (agent0smith):

Again.... KPa is a unit of ***pressure***. T1 is initial temperature and T2 is final temperature V1 is initial volume and V2 is final volume

OpenStudy (anonymous):

I'm confused man

OpenStudy (agent0smith):

er by what...?

OpenStudy (anonymous):

i don't know what to plug into what.

OpenStudy (agent0smith):

Don't plug ANYTHING in until you know what's what. A 20 liter sample of gas is kept under a constant pressure of 100 kPa. The temp is lowered from 360 K to 90 K. What is the new volume read carefully and identify each quantity that you'll need in V1/T1 = V2/T2 T1 is initial temperature and T2 is final temperature V1 is initial volume and V2 is final volume Identify what each one is.

OpenStudy (anonymous):

360 K v1 90k v2 100 kpa t1

OpenStudy (agent0smith):

None of those are correct.

OpenStudy (anonymous):

PLease help me now my brain is turning into liquid

OpenStudy (agent0smith):

You don't seem to know what any of these things are... you need to read your notes or book or something. V1 or V2 is volume. Measured in litres often. T1 or T2 is temperature. Measured in degrees Kelvin often. And again, kPa is a unit of PRESSURE. NOT a measure of temperature or volume.

OpenStudy (anonymous):

t1 90k t2 360k v1 = 100 L

OpenStudy (agent0smith):

Good. Now V2 is your unknown. But make sure you study your notes/book, you can not expect to answer these questions if you have no idea what temperature/volume/pressure are.

OpenStudy (anonymous):

wait my algebra's rusty too. sorry bro im a muscles and no brain/.

OpenStudy (anonymous):

324 cubic liters?

OpenStudy (agent0smith):

\[\Large \frac{ V_1 }{ T_1 } = \frac{ V_2 }{ T2 }\]

OpenStudy (anonymous):

100/30=v2/360

OpenStudy (anonymous):

so...

OpenStudy (anonymous):

100 ____ = v2 /360 30

OpenStudy (anonymous):

36000/30=1200!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (agent0smith):

t1 90k t2 360k v1 = 100 L Take a look at these, then your equation. I don't see a 90 in yours.

OpenStudy (anonymous):

36000/90=400!

OpenStudy (agent0smith):

wait your numbers were wrong. A 20 liter *********V1****** sample of gas is kept under a constant pressure of 100 kPa. The temp is lowered from 360 K *******T1******* to 90 K ********T2******. What is the new volume V2

OpenStudy (agent0smith):

lol btw when i google your profile pic (which obv isn't you haha), google says this about the pic: "Best guess for this image: really gay" hahaha

OpenStudy (anonymous):

oh that cross the line mister! 20 v2 - - 360 90 1800/360 = 5 cubic liters?

OpenStudy (agent0smith):

lol right click your profile pic, then click search google for this image, see for yourself! it's hilarious, the results 5 looks right

OpenStudy (agent0smith):

liters aren't cubic, though. Liters are liters.

OpenStudy (anonymous):

the shooter? what the? lololololololol

OpenStudy (agent0smith):

lol idk this is what google gives me: Best guess for this image: really gay Search Results Who's Really Gay? - YouTube What are some really really gay songs? - Yahoo Answers If Gay Means Stylish, What's Up With That Tacky Rainbow Flag ... etc lol

OpenStudy (anonymous):

-___- your help was appreciated.